OpenStudy (anonymous):

Determine whether the matrix has an inverse. If an inverse exists, find it. matrix: -5 -18 2 7

OpenStudy (anonymous):

inverse exists if determinant is not 0. inverse matrix is given by 1/ad-bc [[d,-b][-c,a]]

OpenStudy (anonymous):

all i see is a bunch of letters..

OpenStudy (anonymous):

\[ \begin{pmatrix} -5 & -18 \\ 2 & -4 \end{pmatrix} \] like that?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

whatever it is, you need to see if the matrix has an inverse before you start to find it do you know how to take the determinant? you need that

OpenStudy (anonymous):

but where did you get the negaticve 4 fromm?

OpenStudy (anonymous):

no i don't.

OpenStudy (anonymous):

\[\begin{pmatrix} -5 & -18 \\ 2 & 7 \end{pmatrix}\]

OpenStudy (anonymous):

is that the one?

OpenStudy (anonymous):

yessss.

OpenStudy (anonymous):

ok determinant of \[\begin{pmatrix} a & b \\ c & d \end{pmatrix}\] is \(ad-bc\) find that first

OpenStudy (anonymous):

okayy im writing it down. you can keep going.

OpenStudy (anonymous):

let me know what you get, then we can continue

OpenStudy (anonymous):

wait im lost... how can you subtract that?

OpenStudy (anonymous):

inverse of \[\begin{pmatrix} a & b \\ c & d \end{pmatrix}\] is easy to find (harder the bigger the matrix) it is \[\frac{1}{ad-bc}\begin{pmatrix} d & -b \\ -c & a \end{pmatrix}\]

OpenStudy (anonymous):

\[a=-5,b=-18, c=2,d=7\] you compute \[(-5)\times 7-(-18\times 2)\]

OpenStudy (anonymous):

okayyy

OpenStudy (anonymous):

i got 1..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which makes it even easier if the determinant is 0 there is no inverse

OpenStudy (anonymous):

to find the inverse use the rule i gave above, but you don't have to divide by the determinant because dividing by 1 is like doing nothing

OpenStudy (anonymous):

ohh, so its not posssible.. it doesn;'t exist?

OpenStudy (anonymous):

no it exists alright

OpenStudy (anonymous):

if the determinant was 0 it would not exist

OpenStudy (anonymous):

but since the determinant is 1 it does exist

OpenStudy (anonymous):

here is the formula \[\frac{1}{ad-bc}\begin{pmatrix} d & -b \\ -c & a \end{pmatrix}\]

OpenStudy (anonymous):

but since \(ad-bc=1\) all you need is \[\begin{pmatrix} d & -b \\ -c & a \end{pmatrix}\]

OpenStudy (anonymous):

so im multiplying everything by one?

OpenStudy (anonymous):

is it clear what you have to do?

OpenStudy (anonymous):

multiplying by one is like doing nothing, so no don't "multiply by one?"

OpenStudy (anonymous):

welll no i don't understand haa. im still lost. im sorrry.

OpenStudy (anonymous):

this is what you need \[\begin{pmatrix} d & -b \\ -c & a \end{pmatrix}\] with \[a=-5,b=-18, c=2,d=7\]

OpenStudy (anonymous):

wait! so your saying its gonna be 5,-18, -2,7? or am i still way off?

OpenStudy (anonymous):

the inverse of a matrix is given by 1 over determinant times adjugate matrix, which for 2x2 matrix is \[\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\]

OpenStudy (anonymous):

so u swap position of a and d and make b and c the opposite sigsn

OpenStudy (anonymous):

sooo idk how to do the matrix box but its gonna be 7 -18 -2 and -5?

OpenStudy (anonymous):